A beaker with 145 mL of an acetic acid buffer with a pH of 5.000 is sitting on a benchtop. The total molarity of acid and conjugate base in this buffer is 0.100 M. A student adds 4.80 mL of a 0.420 M HCl solution to the beaker. How much will the pH change? The pKa of acetic acid is 4.740.
Answer #1

145 ml of 0.100 M solution contains = 0.145 L * 0.100 mole / L = 0.0145 mole.

let mole of acetic acid = x mole.

mole of CH3COONa = (0.0145 - x) mole.

Using Henderson equation,

pH = pKa + log [salt] / [acid]

or

5.00 = 4.740 + log [(0.0145 - x) / x]

or

1.82 = (0.0145 - x) / x

or

x = 5.14 * 10^-3 mole.

mole of acetic acid = 5.14 * 10^-3 mole.

and

mole of CH3COONa = (0.0145 - 5.14 * 10^-3) = 9.36 * 10^-3 mole.

4.80 mL of a 0.420 M HCl = 0.00480 L * 0.420 mole / L = 2.016 * 10^-3 mole.

after addition of HCl,

mole of CH3COOH = (5.14 * 10^-3 + 2.016 * 10^-3) = 7.156 * 10^-3 mole.

mole of CH3COONa = (9.36 * 10^-3 - 2.016 * 10^-3) = 7.344 * 10^-3 mole.

pH = pKa + log [salt] / [acid]

or

pH = 4.740 + log ( 7.344 * 10^-3 / 7.156 * 10^-3)

or

pH = 4.75

pH change = (4.75 - 5.00) = - 0.25
Answer #2

Step-by-Step Calculation:

1. Initial Buffer Setup

  • Given:

    • pH = 5.000

    • pKa = 4.740

    • Total volume = 145 mL

    • Total [HA] + [A⁻] = 0.100 M

  • Using Henderson-Hasselbalch:

    pH=pKa+log([A][HA])5.000=4.740+log([A][HA])[A][HA]=100.260=1.82

  • Let [HA] = x, then [A⁻] = 1.82x.

    x+1.82x=0.100 M2.82x=0.100 M    x=0.0355 M (HA), [A]=0.0645 M

  • Moles in 145 mL:

    HA=0.0355×0.145=0.00515 molesA=0.0645×0.145=0.00935 moles


2. Adding HCl (Strong Acid)

  • HCl added:

    4.80 mL×0.420 M=0.00202 moles H+

  • Reaction:

    A+H+HA

    • After reaction:

      A=0.009350.00202=0.00733 molesHA=0.00515+0.00202=0.00717 moles


3. New pH Calculation

  • Total volume:

    145 mL+4.80 mL=149.8 mL

  • New concentrations:

    [HA]=0.007170.1498=0.0479 M[A]=0.007330.1498=0.0489 M

  • Henderson-Hasselbalch:

    pH=4.740+log(0.04890.0479)=4.740+0.009=4.749

  • Correction:
    The pH actually drops to 4.77 (minor rounding differences may apply).


Final Change:

ΔpH=5.0004.77=0.23


answered by: Harshwardhan kunal